what is the relation between time and charge of electrons

Grade:11

2 Answers

Akachdeep Medhi
5 years ago
We know that electric charges are quantised bodies.this can be given by:
Q=+ or – ?ne (Where e is the charge of an electron and n is any integer)
The relation between electrons(charge),time and electric current can be gven by:
I=Q * t (where I is the electric current and t is the time).
With regards Akash(Class 10)
 
Akachdeep Medhi
5 years ago
The answer should be  I = Q/t instead of I=Q*t. 
I am extremely sorry for that inconvenience.
With regards Akash